# RealGen: Retrieval Augmented Generation for Controllable Traffic Scenarios

### v1

@@ 4.2 The generation pipeline
The retriever enhances the versatility of generation by dividing the process into two stages. In the initial stage, users can choose a set of template scenarios that depict specific conditions. These may include manually annotated scenarios with tags denoting actions like left or right turns, thereby enabling the generation of additional scenarios under similar tags or even a combination thereof. Additionally, templates can encompass critical and interesting scenarios collected from real-world data. Solely relying on these templates for generation could be limited, which is mitigated by incorporating a secondary phase, which employs a KNN approach to fetch high-quality scenarios from a vast and unlabeled database to augment adaptability.
@@ 5.1 Settings and Implementation Details
In RealGen, the implementation of the transformer encoder and multi-head cross attention is based on the PyTorch [42] package. We use Adam [35] as the optimizer for updating the parameters of the auto-encoder and the combiner. To make the calculation of the Wasserstein distance efficient, we use the Sinkhorn distance [13], an entropy regularized approximation of the Wasserstein distance [51], with implementation in the GeomLoss [19] package.
Table 1: Results of realism metrics. Recon-based means using the exact target scenario as input for generation and Retrieval-based means using retrieved scenarios as input for generation. The variance is calculated with 3 seeds.
Category Method mADE mFDE Speed Heading Collision Rate Off-Road Rate
Recon-based AE 0.18$\pm$0.03 0.41$\pm$0.06 0.04$\pm$0.01 0.10$\pm$0.01 0.02$\pm$0.00 0.02$\pm$0.00
@@
Table 2: Accuracy of linear probing
AE Contrastive AE Masked AE RealGen-AE
82.5% 67.1% 86.2% 87.8%
5.2 Baselines
In the experimental section, we evaluate the following reconstruction-based generative models as baselines for comparison. Autoencoder (AE) shares the same behavior encoder, map encoder, and decoder structures as RealGen, serving as the most straightforward baseline for scenario reconstruction. Contrastive AE mirrors the structure of the RealGen autoencoder but omits the initial pose as absolute information. We also consider Masked AE as a self-supervised learning baseline, which has been investigated for trajectory data as outlined in [55, 9, 57, 12]. Given that our method has two phases, we refer to the autoencoder component specifically as RealGen-AE and the complete model as RealGen.
5.3 Evaluation of Behavior Embedding
Linear probing of behavior embedding. Linear probing, a commonly employed method to assess representations in self-supervised learning (SSL), involves training a linear classifier using the derived embeddings [27]. For the purpose of training this classifier, we implemented heuristic rules to assign basic behavioral labels (acceleration, deceleration, stopping, keeping speed, left/right turn) to each agent’s embedding. The outcomes, along with comparisons with baseline models, are presented in Table 2. These findings demonstrate that the embeddings generated by RealGen surpass all baseline models in terms of accuracy.
@@ 5.3 Evaluation of Behavior Embedding
5.4 Evaluation of Retrieval Augmented Generation
Evaluating scene-level controllability and quality of scenario generation is an open problem due to the lack of quantitative metrics. Following most previous works [49, 18, 62], we provide the results of the realism metrics and show the qualitative results of using RealGen for various downstream tasks. More results can be found in the Appendix.
Realism of generated scenario. To evaluate the realism of generated scenarios, we consider the following metrics and show the results in Table 1. We use the maximum mean discrepancy (MMD) [21] to measure the similarity of velocity and heading between the original scenarios and the generated scenarios. We also compare the mean average displacement error (mADE) and the mean final displacement error (mFDE) for the average reconstruction performance. To evaluate scene-level realism, we calculate the scene collision rate and the off-road rate following the metrics defined in [49]. The recon-based generation methods in Table 1 use the behavior of the target scenario as input. For this category, we compare RealGen-AE, only using the encoder and decoder modules, with three baseline methods. Due to the additional contrastive term, RealGen-AE achieves slightly worse performance than AE and Masked AE. However, RealGen is designed for retrieval-based generation, which uses retrieved scenarios rather than the target scenario as input. To fairly compare the generation performance, we design two baselines named AE-KNN and RealGen-AE-KNN, which use KNN to find the most similar behavior embedding to the target scenario and use this embedding as input to the decoder for generation. According to the results, we find that RealGen achieves comparable performance as recon-based generation and outperforms baselines, which indicates the important role of the combiner in fusing the information of the retrieved scenarios.
@@ 5.4 Evaluation of Retrieval Augmented Generation
Crash scenario generation. Beyond the previously mentioned tag, RealGen demonstrates its capacity for in-context learning by generating critical and unseen crash scenarios, divergent from those in the training datasets. This is initiated by manually crafting several crash scenario templates, guided by the scenarios recorded in the NHTSA Crash Report [39]. Subsequently, we generate crash scenarios using existing initial poses and maps of the dataset. Figure 6 illustrates six instances, where the shadowed rectangles denote the initial positions of agents and the red box highlights the point of collision.
@@ 6 Conclusion and Limitation
Supplementary Material
Appendix A Training and Model Details
A.1 Details of model architecture
In addition to the model architecture introduced in Algorithm 1, we also want to provide more details about the spatial-temporal transformer structure. For the projected embedding with shape $[B,M,T,H]$ with batch size $B$, we reshape it to shape $[T,B\times M,H]$ and treat $T$ as the sequence dimension. Then we add a positional embedding to it before putting it into the temporal transformer encoder. After that, we reshape it to $[M,B\times T,H]$ and treat $M$ as the sequence dimension. Then we directly put it into the spatial transformer encoder.
During the decoding process, we repeat the reconstructed embedding $z_{r}$ with shape $[B,M,H]$ along with the temporal dimension to shape $[B,M,T,H]$. Then we follow the same process as the encoding process for spatial temporal decoding with a positional embedding added before the temporal transformer encoder.
A.2 Details of metrics
We provide details about the metrics we use in the experiment section.
Maximum Mean Discrepancy (MMD) is a statistical measure used to test the similarity of two distributions, denoted as $P$ and $Q$, by calculating the distance between the means of these distributions in a feature space defined by a kernel function. In our implementation, we use a Gaussian kernel $k(x,y)$ (also known as Radial Basis Function kernel). Assume we have samples $X=\{x_{1},x_{2},...,x_{m}\}$ from distribution $P$ and $Y=\{y_{1},y_{2},...,y_{n}\}$ from distribution $Q$, we calculate the MMD statistic as:
$\begin{split}&\text{MMD}^{2}(P,Q)=\\ &\frac{\sum_{i,j}k(x_{i},x_{j})}{m^{2}}+\frac{\sum_{i,j}k(y_{i},y_{j})}{n^{2}}+\frac{2\sum_{i,j}k(x_{i},y_{j})}{mn},\end{split}$ (7)
where $m$ and $n$ are the number of samples from $P$ and $Q$ respectively. A larger MMD value indicates a greater discrepancy between the two distributions.
Mean Average Displacement Error (mADE) is a metric commonly used in the field of trajectory prediction, which quantifies the accuracy of predicted trajectories by comparing them with the actual trajectories. Assume the predicted trajectory is $\hat{\tau}_{i}=\{\hat{p}^{i}_{1},\hat{p}^{i}_{2},...,\hat{p}^{i}_{T}\}$ and the ground truth trajectory is $\tau_{i}=\{{p}^{i}_{1},{p}^{i}_{2},...,{p}^{i}_{T}\}$ for agent $i$ with trajectory length $T$, we calculate mADE with
$\text{mADE}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{T}\sum_{t=1}^{T}\|\hat{p}^{i}_{t}-p^{i}_{t}\|\right),$ (8)
where $N$ is the number of agents in the validation dataset.
Mean Final Displacement Error (mFDE) is another important metric used in trajectory prediction, similar to mADE, but with a specific focus on the prediction accuracy at the final time step of the trajectories. With the same notation defined above, we calculate mFDE with
$\text{mFDE}=\frac{1}{N}\sum_{i=1}^{N}\|\hat{p}^{i}_{T}-p^{i}_{T}\|.$ (9)
Scene Collision Rate is used to calculate the overlap between two vehicles for every timesteps in every trajectories. Specifically, a collision is said to occur between two vehicles if the Intersection Over Union (IOU) of their representing rectangles exceeds the threshold $\theta$:
$\text{Collision}(R_{1},R_{2})=\left\{\begin{array}[]{l}1,\text{if IOU}(R_{1},R_{2})>\theta\\ 0,\text{otherwise}\end{array}\right.$ (10)
where the threshold $\theta$ is set to 0.1 and IOU is defined as:
$\text{IOU}(R_{1},R_{2})=\frac{\text{Area of Intersection ($R_{1}$, $R_{2}$)}}{\text{Area of Union ($R_{1}$, $R_{2}$)}}.$ (11)
Therefore, the scene collision rate is defined as:
Scene Collision $\text{Scene Collision Rate}=\frac{1}{N}\sum_{i=1}^{N}\text{Collision}(R_{1},R_{2}).$ (12)
Off-Road Rate is used to calculate whether the trajectory is in the drivable area. We first calculate the indicator function of off-road with
[…]
Then we calculate the rate with
[…]
A.3 Hyper-parameters
We list all important hyper-parameters we used during the training of the autonecoder and the combiner and the generation process in Table 3.
A.4 Computational resources
For all experiments, we use a server with AMD EPYC 7763 64-Core CPU, 256GB memory, and 4$\times$×RTX A6000. The training of the autoencoder consumes 7.5GB memory and requires 8 hours on 1$\times$×RTX A6000. The training of the combiner consumes 2.1GB memory and requires 3 hours on 1$\times$×RTX A6000.
Table 3: Hyper-parameters
Parameter Name Parameter Notation Parameter Value
Max timestep $T$ 17
Max agent number $M$ 11
Trajectory frequency 2Hz
Trajectory dimension meaning [x, y, v, cos(h), sin(h)]
Map dimension meaning [x, y, cos(h), sin(h)]
KNN Retrieve number $K$ 5
Number of epochs (Autoencoder) 300
Batch size (Autoencoder) 64
Initial learning rate (Autoencoder) 0.0008
Learning rate scheduler (Autoencoder) MultiStepLR
Scheduler parameter (Autoencoder) [20, 40, 60, 80, 100, 200]
Gradient clip 5.0
Weight of InfoNCE loss $\lambda$ 0.1
Temperature of InfoNCE $\tau$ 0.1
Moments of Wasserstein distance $p$ 2
Blur of Wasserstein distance 0.05
Number of epochs (Combiner) 500
Batch size (Combiner) 64
Initial learning rate (Combiner) 0.001
Learning rate scheduler (Combiner) MultiStepLR
Scheduler parameter (Combiner) [20, 40, 60, 80, 100, 200]
Radius for searching lanes 100 m
Number of points in lane 20
Max number of lane 100
Threshold for filtering static agent 3 m
Latent embedding dimension dim($z_{b}$), dim($z_{i}$), dim($z_{m}$) 256
number of attention head 16
transformer feedforward dimension 512
dropout rate 0.1
# Behavior encoder layer $L_{e}$ 2
# Behavior Spatial TransEnc layer 2
# Behavior Temporal TransEnc layer 1
Behavior projection MLP [5, 512]
Initial pose projection MLP [5, 512]
Map projection MLP [4, 512]
Map encoding MLP [512, ReLU, Dropout, 512]
# Decoder layer number $L_{d}$ 2
# Decoder Spatial TransEnc layer 2
# Decoder Temporal TransEnc layer 1
Decoder projection MLP [512, ReLU, 512, ReLU, 5]
Appendix B More Experiment Results
We provide more experiment results in this section to demonstrate the performance of our method.
B.1 Visualization of similar scenarios
In Figure 7, we show more examples of scenarios similar and dissimilar to the query scenario.
Figure 7: More examples of qualitative evaluation of similar and dissimilar scenarios calculated by our scenario embedding.

### v2

@@ 4.2 The generation pipeline
The retriever enhances the versatility of generation by dividing the process into two stages. In the initial stage, users can choose a set of template scenarios that depict specific conditions. These may include manually annotated scenarios with tags denoting actions like left or right turns, enabling the generation of additional scenarios under similar tags or even a combination thereof. In addition, templates can include critical and interesting scenarios collected from real-world data. Solely relying on these templates for generation could be limited, which is mitigated by incorporating a secondary phase, which employs a KNN approach to fetch high-quality scenarios from a vast and unlabeled database to augment adaptability.
@@ 5.1 Settings and Implementation Details
To train RealGen, we use Adam [36] as the optimizer to update the parameters of the autoencoder and the combiner. To make the calculation of the Wasserstein distance efficient, we use the Sinkhorn distance [13], an entropy regularized approximation of the Wasserstein distance [52], with implementation in the GeomLoss [19] package.
5.2 Baselines
In the experimental section, we evaluate the following reconstruction-based generative models as baselines for comparison. Autoencoder (AE) shares the same behavior encoder, map encoder, and decoder structures as RealGen, serving as the most straightforward baseline for scenario reconstruction. Contrastive AE mirrors the structure of the RealGen autoencoder but omits the initial pose as absolute information. Masked AE is a self-supervised learning baseline, which has been investigated for trajectory data as described in [56, 9, 58, 12]. To evaluate the controllable generation capability, we select a state-of-the-art modelLCTGen [49] as the baseline, which takes a high-level agent representation $z$ and a vector map $m$ as input. We also consider LCTGen w/o $z$, a variant model proposed in the original paper to show the reconstruction results without using the agent information. Given that our method has two phases, we refer to the autoencoder component specifically as RealGen-AE and the complete model as RealGen.
5.3 Evaluation of Behavior Embedding
Table 1: Accuracy of linear probing
AE Contrastive AE Masked AE RealGen-AE
82.5% 67.1% 86.2% 87.8%
Linear probing of behavior embedding. Linear probing, a commonly employed method to assess representations in self-supervised learning (SSL), involves training a linear classifier using the derived embeddings [27]. To train this classifier, we implemented heuristic rules to assign basic behavioral labels (acceleration, deceleration, stopping, keeping speed, left/right turn) to each agent’s embedding. The outcomes, along with comparisons with baseline models, are presented in Table 1. These findings demonstrate that the embeddings generated by RealGen surpass all baseline models in terms of accuracy.
5.4 Evaluation of Retrieval Augmented Generation
Evaluating scene-level controllability and quality of scenario generation is an open problem due to the lack of quantitative metrics. Following most previous work [49, 18, 64], we provide the results of the realism metrics and show the qualitative results of using RealGen for various downstream tasks. More results can be found in the appendix.
Table 2: Results of realism metrics. Recon.-based and Retrieval-based means using the target scenario and retrieved scenarios as input for generation, respectively.
Category Method mADE mFDE Speed Heading SCR ORR
Recon. -based AE 0.18$\pm$0.03 0.41$\pm$0.06 0.04$\pm$0.01 0.10$\pm$0.01 0.02$\pm$0.00 0.02$\pm$0.00
@@ 5.4 Evaluation of Retrieval Augmented Generation
Realism of generated scenario. To evaluate the realism of generated scenarios, we consider the following metrics and show the results in Table 2. We use the maximum mean discrepancy (MMD) [22] to measure the similarity in velocity and direction between the original scenarios and the generated scenarios. We also compare the mean average displacement error (mADE) and the mean final displacement error (mFDE) for the average reconstruction performance. To evaluate scene-level realism, we calculate the scene collision rate (SCR) and the off-road rate (ORR) following the metrics defined in [49]. The recon-based generation methods in Table 2 use the behavior of the target scenario as input. For this category, we compare RealGen-AE, only using the encoder and decoder modules, with three baseline methods. Due to the additional contrastive term, RealGen-AE achieves slightly worse performance than AE and Masked AE. However, RealGen is designed for retrieval-based generation, which uses retrieved scenarios rather than the target scenario as input. To fairly compare the generation performance, we design two baselines named AE-KNN and RealGen-AE-KNN, which use KNN to find the most similar behavior embedding to the target scenario and use this embedding as input to the decoder for generation. According to the results, we find that RealGen achieves comparable performance as recon-based generation and outperforms baselines, which indicates the important role of the combiner in fusing the information of the retrieved scenarios.
@@
Table 3: Safety-critical scenario generation.
Method Realgen-AE-R RealGen-R RealGen
Collision Rate ($\uparrow$) 0.92 0.83 0.59
Safety-critical scenario generation. Beyond the tag mentioned above, RealGen demonstrates its capacity for in-context learning by generating critical and unseen crash scenarios, divergent from those of the training datasets. This is initiated by manually creating several crash scenario templates, guided by the scenarios recorded in the NHTSA Crash Report [40]. Subsequently, we generate crash scenarios using existing initial poses and maps of the dataset. Figure 6 illustrates six instances in which the shadowed rectangles denote the initial positions of agents and the red box highlights the point of collision. To quantitatively assess the performance of safety-critical scenarios, we compare RealGen with two baselines. RealGen-AE-R random samples behavior embeddings in the Realgen-AE model, and RealGen-R random retrieves behavior embeddings. According to the results in Table 3, we find that the scenarios generated by RealGen using crash scenarios as templates achieve the highest collision rate, which means that RealGen has more efficiency.
Human evaluation of controllability. As there is no automatic way to evaluate controllability, we follow the protocol in [49] to perform A/B testing using human evaluation. We report the ratio of our method preferred as well as the absolute score (0-5) of both our method and the baseline. In Table 4, we find that scenarios generated by RealGen are highly preferred in most categories. The absolute score of RealGen is also much higher than that of LCTGen.
Table 4: Results of human evaluation of controllability. (Details in Appendix LABEL:app:human)
Category Left-Turn Right-Turn Left-Lane-Change Right-Lane-Change Straight
[…]
RealGen Score (0-5) 4.27$\pm$1.05 4.27$\pm$0.69 3.96$\pm$1.94 4.17$\pm$1.93 3.94$\pm$2.27
LCTGen Score (0-5) 2.15$\pm$1.44 2.08$\pm$1.19 2.42$\pm$1.95 2.0$\pm$1.96 2.14$\pm$2.31
Table 5: Downstream task evaluation.
Method Original Random Aug. RealGen Aug.
mADE ($\downarrow$) 3.544 2.920 2.309
collision rate ($\downarrow$) 0.049 0.037 0.018
Downstream task evaluation. A direct downstream task of our method is to use the generated data to augment the training dataset of trajectory prediction models. We use Autobots [21] as a predictor and report the results trained on different datasets in Table 5. Original means using the original data in nuScenes [5], Random Aug. means augmenting the original dataset with Gaussian noise, and RealGen Aug. means augmenting the original dataset with scenarios from RealGen. We observe that the model trained with the RealGen dataset achieves the lowest mADE and collision rate.
@@ 6 Conclusion and Limitation
Acknowledgments
Wenhao Ding contributed to this paper while being an intern at NVIDIA Research. Ding Zhao was partially supported by the National Science Foundation under grant CNS-2047454.